Night Report 2016-04-03

Observing: Joel Johansson, Mark Magee, Ira Bar

Support Team: QUB

TAT: Lluís Galbany, Joe Lyman

Weather conditions

End of night: cloudy, humidity ~64%, wind ~12m/s, seeing ~1.5"


Start of night: thick clouds, humidity ~63%, wind ~14 m/s

Middle of night: cloudy, humidity ~60%, wind ~13m/s, seeing ~1.2"


Evening: PESSTO_bias, PESSTO_calib

Morning: Dome flats


01:33 UT: Dome opened, pointing restriction DEC <-20

02:04 UT: Follow up: Gaia16aec, gr13, 3x2700s, 1" slit*

04:25 UT: Dome closed due to clouds

06:02 UT: Dome opened, pointing restriction

06:10 UT: Image analysis

06:27 UT: Follow up: SN2016aj, ri, 8x100s, seeing ~1.6"

07:14 UT: Standard: LTT7379, gr11+13+16, 1" + 1.5", 60s each

07:32 UT: Follow up: SN2016adj, gr11 2000s, gr16 400s, 1.5" slit

08:40 UT: Follow up: SN2016adj, BVRi imaging, 10 mins

08:58 UT: Classification: AT2016blj, gr13, 600s, 1.5" slit

09:40 UT: Classification: LSQ16ahs, gr13, 600s, 1.5" slit

09:50 UT: Dome closed

*Second exposure aborted with ~100s remaining due to rotator error. 3rd exposure aborted with ~13 min remaining due to clouds/dome closed.